Polyester division of Quintex Ltd has forecast a net profit before tax of R3 million per annum for the next five years, based on net capital employed of R10 million. Plant replacement over this period is expected to be equal to the annual depreciation each year. These figures compare well with the group’s required rate of return of 20% before tax.

Polyester’s management is currently considering a substantial expansion of its manufacturing capacity to cope with the forecast demands of a new customer.

• The customer is prepared to offer a five year contract providing Polyester with annual sales of R2 million.

• In order to meet this contract, a total additional capital outlay of R2 million is envisaged, being R1.5 million of new fixed assets plus R0.5 million of working capital. The plant life is expected to be 5 years with zero scrap value.

• Operating costs for the contract are estimated to be R1.35 million per annum, excluding depreciation.

• This is considered to be a low-risk venture as the contract would be firm for 5 years and the manufacturing processes are well understood within Polyester.

• The consequences of income tax on the proposal may be ignored.

Required:

Calculate the impact of accepting the contract on Polyester division’s:

(a) Return on investment (ROI) for each of the 5 years

(b) Residual income (RI), using 20% imputed interest rate, using average capital employed per annum, and say, with reasons, for each method whether or not it would be attractive to Polyester division’s management.

Performance Evaluation (Management Accounting)
Answer #1
W.N
Annual depreciation = R1.5 million / 5 years = R0.3 million per annum
Year Net value of fixed assets at end of year Working capital Total capital employed at end of year Average capital employed during year
R million R million R million R million
1 1.2 0.5 1.7 1.85
2 0.9 0.5 1.4 1.55
3 0.6 0.5 1.1 1.25
4 0.3 0.5 0.8 0.95
5 0 0.5 0.5 0.65
Annual net profit R million
Sales 2
Operating costs 1.35
Depreciation (R1.5 million / 5 years ) 0.3
Net profit per annum 0.35
a and b
Year ROI% = Net Profit per annum/Average capital employed during year ROI Imputed interest = Average capital employed during year x 20% Imputed Interest Net profit Residual income = Net Profit - Imputed Interest Residual Income
% R million R million R million
1 0.35 / 1.85 = 18.92% 1.85 x 20% = 0.37 0.35 .35 - .37 -0.02
2 0.35 / 1.55 = 22.58% 1.55 x 20% = 0.31 0.35 .35 - .31 0.04
3 0.35 / 1.25 = 28.00% 1.25 x 20% = 0.25 0.35 .35 - .25 0.1
4 0.35 / 0.95 = 36.84% 0.95 x 20% = 0.19 0.35 .35 - .19 0.16
5 0.35 / 0.65 = 53.85% 0.65 x 20% = 0.13 0.35 .35 - .13 0.22
Current Average 32.04%
If Polyester’s performance is measured using ROI, then management will be reluctant to accept the contract because the ROI in years 1 – 3 is lower than the current average of 32%.
If Polyester’s performance is measured using RI then management would probably still be reluctant to accept the contract because of the negative result in year 1.